OpenStudy (anonymous):

Solve square root of x plus 2 end root minus 2 equals x

OpenStudy (anonymous):

type this the right way, use the equation bottom under the text for the symbols because i don't understand what yo are trying to say

OpenStudy (anonymous):

square both sides to get ride of the square root and then use the quadratic forumla to solve for x

OpenStudy (anonymous):

\[\sqrt{x-2}-2=x\]

OpenStudy (anonymous):

@chinenyeogueri This is their first question, I'm pretty sure they haven't seen the equation tool until after they typed their question since you can't use it for typing a question, but you can see it here

OpenStudy (anonymous):

\[\Large (\color{red}a+\color{blue}b)^2=\color{red}a^2+\color{green}{2ab}+\color{blue}b^2\] You're going to need to use this later on in this question

OpenStudy (anonymous):

Okay i think i got it, Thanks so much!

OpenStudy (anonymous):

Anytime. Just come back if you get lost.

OpenStudy (anonymous):

\[\sqrt{x+2}-2=x \]\[\{x=-2,x=-1\} \]

OpenStudy (mathstudent55):

Start by adding 2 to both sides to isolate the square root. Then square both sides. This gets rid of the square root. Then combine like terms and solve the quadratic equation. Finally, check every answer in the original equation because squaring both sides can introduce extraneous equations.

OpenStudy (yanasidlinskiy):

@jaredbarnes17 , \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)